OpenStudy (anonymous):

What is the value of x? 4(6 + 7) = 4 • 6 + x • 7

TheSmartOne (thesmartone):

do you know the distributive property?

OpenStudy (iwillrektyou):

To solve for x, you first have to solve for the rest/combine like terms.

TheSmartOne (thesmartone):

knowing the distributive property can save you time ;p

TheSmartOne (thesmartone):

\(\sf a(b+c) = a \cdot b + a\cdot c\)

OpenStudy (hockeychick23):

4(6 + 7) = 4 • 6 + x • 7 First we distribute the 4 on the left hand side--> 4*6+4*7= ?

TheSmartOne (thesmartone):

-.-